OpenStudy (anonymous):

Solve the triangle a = 12, b = 14, c = 21

OpenStudy (anonymous):

Apply law of Cosines to find the angle!

OpenStudy (anonymous):

how do you do that?

OpenStudy (anonymous):

Post the formula, I'll show you :)

OpenStudy (anonymous):

c^2 = a^2 + b^2 – 2ab cos C and a^2 = b^2 + c^2 – 2bc cos A and b^2 = a^2 + c^2 – 2ac cos B.

OpenStudy (anonymous):

a² = b² + c² – 2bc cos A ->cosA = ( b² + c² - a² ) / 2bc

OpenStudy (anonymous):

how did you change that?

OpenStudy (anonymous):

We find the angle, so we need to switch around :0

OpenStudy (anonymous):

okay so a would be 84 degrees?

OpenStudy (anonymous):

How???

OpenStudy (anonymous):

Don't just give me your result, because I don't have the power to read your mind :P

OpenStudy (anonymous):

(14^2+21^2-12^2)=493 493/2bc 2bc=2*14*21=588 493/588=0.838 since we are finding degrees that would be 84 degrees right? and sorry lol

OpenStudy (anonymous):

My calculator shows 33° !

OpenStudy (anonymous):

How did you do that? im confused i used my graphing calculator......

OpenStudy (anonymous):

A = arccos (.838) =

OpenStudy (anonymous):

what is arccos?

OpenStudy (anonymous):

oh nvm i found it...

OpenStudy (anonymous):

inverse of cos to find the angle :0

OpenStudy (anonymous):

okay so the how would we find B?

OpenStudy (anonymous):

Now, use sines law to find the second angle!

OpenStudy (anonymous):

b^2 = a^2 + c^2 – 2ac cos B. or cos B=(a^2+c^2-b^2)/2ac right?

OpenStudy (anonymous):

oh sin A/a = sin B/b = sin C/c

OpenStudy (anonymous):

yes, SinB = b sinA/a

OpenStudy (anonymous):

Sin B= 14*sin(33 Degrees)/12 Sin B=14*.045 Sin B= .635 B= 39 degrees?

OpenStudy (anonymous):

Yup, I can see you make great progress =)

OpenStudy (anonymous):

And C would be Law of Sines too?

OpenStudy (anonymous):

No, C = 180° - ( A + B) = ...

OpenStudy (anonymous):

oh okay you just add the two angle and subtract because it is the only angle left in the triangle... So... C=180 degrees -(33+39) C= 180 degrees - 72 degrees C= 108 Degrees

OpenStudy (anonymous):

I have three more solve the triangle problems but they are different... will you help... PLZ

OpenStudy (anonymous):

C= 107.39°

OpenStudy (anonymous):

Just open the new post, I'll follow up :)

OpenStudy (anonymous):

Okay THANK YOU SO SO MUCH!!!!!!!!!!!!!!!!!!!!